a brace for a shelf has the shape of a right triangle. its hypotenuse is 18 inches and the legs are equal in length. How long are the legs of the triangle
Answer by rfer(16322)

a^2+b^2=c^2
18^2=324
324/2=162
sqrt 162=12.73 in